# Dynamics-Aware Gaussian Splatting Streaming Towards Fast On-the-Fly Training for 4D Reconstruction

### v1

@@ Abstract
The recent development of 3D Gaussian Splatting (3DGS) has led to great interest in 4D dynamic spatial reconstruction from multi-view visual inputs. While existing approaches mainly rely on processing full-length multi-view videos for 4D reconstruction, there has been limited exploration of iterative online reconstruction methods that enable on-the-fly training and per-frame streaming. Current 3DGS-based streaming methods treat the Gaussian primitives uniformly and constantly renew the densified Gaussians, thereby overlooking the difference between dynamic and static features and also neglecting the temporal continuity in the scene. To address these limitations, we propose a novel three-stage pipeline for iterative streamable 4D dynamic spatial reconstruction. Our pipeline comprises a selective inheritance stage to preserve temporal continuity, a dynamics-aware shift stage for distinguishing dynamic and static primitives and optimizing their movements, and an error-guided densification stage to accommodate emerging objects. Our method achieves state-of-the-art performance in online 4D reconstruction, demonstrating a 20% improvement in on-the-fly training speed, superior representation quality, and real-time rendering capability. Project page: https://www.liuzhening.top/DASS
Figure 1: Performance of 4D dynamic spatial reconstruction methods in terms of per-frame training time and reconstruction quality (PSNR) on the N3DV dataset [26]. Online methods are represented by circles, while offline methods are indicated by triangles. The size of each marker is proportional to the rendering speed (FPS). Our method achieves the fastest online training speed, superior reconstruction quality, and real-time rendering performance.
@@ 1 Introduction
In recent years, Neural Radiance Field (NeRF) [33] has emerged as a promising approach for spatial representation and reconstruction. NeRFs optimize neural networks to estimate color and density based on spatial position and viewpoint for 3D reconstruction using captured multi-view inputs. The extensions of NeRFs to dynamic scene reconstruction have demonstrated significant effectiveness [26, 35, 37, 14, 15], yielding photo-realistic novel view synthesis results. However, the efficiency of NeRF-based methods is severely hindered by their low rendering speed due to the dense queries of neural networks. To address this issue, 3D Gaussian Splatting (3DGS) [21] has been proposed as a solution to provide high-quality reconstruction and real-time rendering capabilities, leveraging its flexible point-based primitive design and tile-based differentiable rasterization. Subsequent research efforts [19, 52, 27, 47, 3] have been dedicated to applying Gaussian Splatting for 4D dynamic reconstruction, with representative works integrating the time dimension into each Gaussian primitive [52] or learning spatio-temporal deformations [27, 47].
@@
Despite these advancements, most NeRF-based and 3DGS-based methods for dynamic spatial reconstruction rely on full-length multi-view videos, i.e., non-causal inputs. This reliance overlooks applications such as live streaming, where only per-frame causal inputs are available and on-the-fly training is required. This scenario is formalized as iteratively reconstructing 3D space at the current frame based on previous reconstruction caches and current multi-view inputs. The key challenges in this context are two-fold: (i) how to model temporal variations between frames in 3D space and (2) how to facilitate the optimization convergence from the previous frame to the current one. Critical metrics in this scenario include both the quality of novel view synthesis and streaming time efficiency.
@@
One intuitive solution is to directly optimize a new set of 3DGS primitives for each frame. However, tuning and storing all 3DGS parameters for each frame results in significant time costs and storage overhead. A representative baseline, 3DGStream [38], efficiently optimizes the transformation of Gaussian positions and rotation quaternions, and adaptively densifies a small number of new Gaussians. Although this method achieves fast and high-quality results, it overlooks the difference between inherent dynamics and statics in the scene, instead treating the whole scene uniformly. When modeling movements in the space, dynamic and static components showcase different deformation characteristics. For instance, moving objects, such as humans or animals, may display substantial dynamics, with the Gaussian properties like position experiencing significant offsets. In contrast, static background and stationary objects show minimal movement, where Gaussians remain unchanged or undergo slight jitters. Besides, in most natural scenes, only a small subset of Gaussian primitives corresponds to dynamic areas. Consequently, uniformly modeling the transformation of all Gaussians results in a sub-optimal solution. Moreover, renewing the added Gaussian primitives for each frame fails to fully exploit the temporal continuity.
@@
Based on these insights, we propose a dynamics-aware 3DGS streaming paradigm for on-the-fly 4D reconstruction, termed DASS, where the optimization of each frame comprises three stages: inheritance, shift, and densification. Specifically, considering the temporal continuity, the newly added Gaussians in the previous frame are likely to persist in subsequent frames. Therefore, instead of renewing these added Gaussians for each frame and optimizing them from scratch, we propose a selective inheritance mechanism to adaptively include a portion of the added Gaussians from the previous frame using a learnable selection mask. Then, in the shift stage, we employ 2D dynamics-related prior optical flow [48, 49, 45] and Gaussian segmentation [53, 36] to calculate a per-Gaussian dynamics mask. Subsequently, we assign grid-based layers to learn the offsets of dynamic and static Gaussians with different representation complexities. In the densification stage, apart from the Gaussian offsets that present the deformations of existing objects, new Gaussian primitives are introduced to accommodate newly emerging objects. In this stage, both positional gradients and error maps from the shift stage serve as criteria for identifying regions that require densification. The inheritance stage in the subsequent frame will process these added Gaussians, thereby mitigating errors in the shift stage and reducing the optimization burden in the densification stage. This three-stage pipeline effectively captures dynamic spatial components and exploits the temporal correlation, providing fast on-the-fly training and high-fidelity streaming. Our main contributions are summarized as follows:
@@
  • We propose a novel three-stage pipeline for 4D dynamic spatial reconstruction that supports on-the-fly training and per-frame streaming. Our method builds on causal inputs and eliminates the need for full-length multi-view videos, thereby enhancing the practicability.
@@
  • Our approach seamlessly integrates the three stages to optimize the reconstruction quality. By selectively inheriting newly introduced Gaussians from the preceding frame, effectively distinguishing dynamic primitives to allocate optimization emphasis, and enhancing areas with weak reconstruction using gradient information and optimization errors, our method ensures high-fidelity dynamic spatial reconstruction.
@@
  • Extensive experiments demonstrate the superiority of our method in multiple aspects, including a 20% improvement in online training speed, superior reconstruction quality, and real-time rendering capability.
@@ 2.1 Neural Static Scene Representation
In recent years, reconstructing 3D representations from 2D plane visual inputs has experienced significant advancements, driven by the development of NeRF [33]. NeRF-based methods represent spatial scenes by optimizing multi-layer perceptrons (MLPs) and generate novel views through volume rendering [18]. Subsequent research has enhanced both the training and rendering efficiency through grid-based designs [34, 50, 13, 14]. Nonetheless, NeRF-based approaches typically require dense ray tracing and struggle to fulfill high-speed rendering. Recently, 3DGS [21] has emerged to address these limitations by utilizing explicit unstructured scene representation while preserving point-based differentiable splatting rendering [57]. This approach achieves real-time rendering speed and photo-realistic quality. Based on these advancements, subsequent studies have further enhanced the representation efficiency [31, 8, 30, 11, 24, 55], developed feed-forward reconstruction models [56, 29, 6, 17], and expanded applications in understanding and editing [53, 7, 36, 44].
@@ 2.2 Neural Dynamic Scene Reconstruction
Extending static scene representation to dynamic FVV reconstruction remains a significant challenge, primarily due to the difficulties in modeling temporal correlations and variations in 3D space. To address this issue, several studies have extended NeRF to incorporate spatio-temporal structures [25, 5, 37, 14, 42], facilitating dynamic space reconstruction. Similarly, dynamic scene reconstruction methods based on Gaussian Splatting have been proposed, which can be categorized into deformation-based methods, 4D primitive-based methods, and iterative streaming methods. Deformation-based methods [47, 32, 51, 3] maintain 3D Gaussian representations and optimize a neural deformable field to capture temporal variations. 4D primitive-based methods [52, 27, 54] augment Gaussian primitives by integrating the temporal axis as an intrinsic property, thereby directly learning the spatio-temporal distributions. Note that both categories of methods rely on full-length multi-view videos for training, which limits their ability to support on-the-fly training and per-frame streaming. In contrast, iterative streaming methods build upon previously converged representations and perform per-frame optimization for each frame’s multi-view inputs, thereby addressing the above limitations. However, this area remains largely underexplored. Our work aims to develop an iterative streaming pipeline that accelerates per-frame optimization convergence and enhances the reconstruction quality.
The work most closely related to ours is 3DGStream [38]. It proposes a two-stage optimization process for each frame, where the first stage optimizes a grid-based MLP [34] for Gaussian property transformation and the second stage wisely adds new Gaussians based on positional gradients. However, 3DGStream does not consider the intrinsic dynamic and static features in the scene and erases the added primitives in each frame, which neglects the temporal consistency. We have also noticed several concurrent studies on iterative 4D reconstruction, although their focuses differ from ours. For instance, S4D [16] directly manipulates 3D control points to guide the movements, but it requires full optimization that incurs significant time costs on key frames. V3 [43] aims to facilitate streaming free-viewpoint videos on mobile devices, with a particular emphasis on the reduction of the per-frame streaming storage overhead. Conversely, our work identifies the per-frame convergence speed as the critical bottleneck. Moreover, while V3 decomposes motion and appearance attributes for efficient reconstruction, it primarily applies to scenarios with moving objects against vacant background and does not accommodate emerging objects. In addition, SwinGS [28] concentrates on maintaining a consistently stable stream data volume over time and reducing storage overhead through the use of Markov Chain Monte Carlo [22] and a window-based design. Another work [39] considers the bit allocation in a system-level implementation. In contrast, our work focuses on achieving fast convergence in per-frame optimization while maintaining high reconstruction quality, which is orthogonal to the contributions of aforementioned studies.
@@
Figure 2: Overview of our proposed DASS framework. The selective inheritance stage (Green) exploits the temporal continuity and adaptively preserves Gaussians from the previous frame. The dynamics-aware shift stage (Blue) distinguishes the dynamic and static elements and optimizes the deformations. The error-guided densification stage (Yellow) detects and densifies the areas with weak reconstruction based on positional gradients and distortions. Variables highlighted in red represent learnable parameters in each stage, whose training is significantly lightweight compared to tuning all Gaussian parameters.
@@ 3.1 Overview
Our method, referred to as DASS, achieves 4D dynamic spatial reconstruction in an iterative manner, enabling on-the-fly training and per-frame streaming. At each timestep $t$, the optimization begins with the representation of the previous timestep, denoted by $\bm{\mathcal{G}}_{t-1}$, to derive the current representation $\bm{\mathcal{G}}_{t}$, based on the current multi-view inputs from $C$ viewpoints. For each scene, the pipeline starts from an initial set of 3D Gaussians at frame 0. This initial representation $\bm{\mathcal{G}}_{0}$ consistently serves as a base representation $\bm{\mathcal{G}}^{\textrm{base}}$ with a constant number of primitives $N$ across subsequent frames. Moreover, each timestep incorporates a varying yet limited number of densified Gaussians, denoted by $\bm{\mathcal{G}}^{\textrm{densi}}$, to model the unique per-frame objects. A subset of these densified Gaussians is inherited to the next timestep as $\bm{\mathcal{G}}^{\textrm{inher}}$.
@@
Fig. 2 depicts an overview of our DASS framework, where the optimization for each frame comprises three stages: inheritance, shift, and densification. Unlike previous methods that regenerate densified Gaussians at each timestep without considering temporal continuity, our approach introduces a selective inheritance stage (Sec. 3.2) at the beginning of optimization for each timestep. During this stage, the densified Gaussians from the preceding timestep, $\bm{\mathcal{G}}^{\textrm{densi}}_{t-1}$, are selectively inherited using a per-Gaussian learnable mask. This mask effectively eliminates redundant Gaussians while preserving essential ones, transforming the set from $\{\bm{\mathcal{G}}^{\textrm{base}}_{t-1},\bm{\mathcal{G}}^{\textrm{densi}}_{t-1}\}$ to $\{\bm{\mathcal{G}}^{\textrm{base}}_{t-1},\bm{\mathcal{G}}^{\textrm{inher}}_{t-1}\}$. Then, the shift stage (Sec. 3.3) processes the base Gaussians $\bm{\mathcal{G}}^{\textrm{base}}_{t-1}$ and the inherited densified Gaussians $\bm{\mathcal{G}}^{\textrm{inher}}_{t-1}$. In this stage, a Gaussian-level dynamics mask is estimated to categorize the Gaussians into dynamic and static groups. These categories are then processed by two hash-encoding MLPs that operate at different complexities. This shift stage efficiently manages the movements and rotations of Gaussian primitives, accommodating changes from the previous timestep and updating the representation as $\{\bm{\mathcal{G}}^{\textrm{base}}_{t-1},\bm{\mathcal{G}}^{\textrm{inher}}_{t-1}\}\to\{\bm{\mathcal{G}}^{\textrm{base}}_{t},\bm{\mathcal{G}}^{\textrm{inher}}_{t}\}$. In the densification stage (Sec. 3.4), emerging objects, such as coffee being poured from a cup and flames from an oven, are identified by analyzing positional gradients and distortions resulted from the shift stage. Then, this stage densifies additional Gaussian primitives to accurately represent these emerging objects and yields $\{\bm{\mathcal{G}}^{\textrm{base}}_{t},\bm{\mathcal{G}}^{\textrm{inher}}_{t}\}\to\{\bm{\mathcal{G}}^{\textrm{base}}_{t},\bm{\mathcal{G}}^{\textrm{densi}}_{t}\}$. Our proposed design facilitates fast per-frame convergence and produces high-fidelity view synthesis results. The details of each stage are elaborated in the following subsections.

### v2

# Dynamics-Aware Gaussian Splatting Streaming Towards Fast On-the-Fly 4D Reconstruction

@@ Abstract
The recent development of 3D Gaussian Splatting (3DGS) has led to great interest in 4D dynamic spatial reconstruction. Existing approaches mainly rely on full-length multi-view videos, while there has been limited exploration of online reconstruction methods that enable on-the-fly training and per-timestep streaming. Current 3DGS-based streaming methods treat the Gaussian primitives uniformly and constantly renew the densified Gaussians, thereby overlooking the difference between dynamic and static features as well as neglecting the temporal continuity in the scene. To address these limitations, we propose a novel three-stage pipeline for iterative streamable 4D dynamic spatial reconstruction. Our pipeline comprises a selective inheritance stage to preserve temporal continuity, a dynamics-aware shift stage to distinguish dynamic and static primitives and optimize their movements, and an error-guided densification stage to accommodate emerging objects. Our method achieves state-of-the-art performance in online 4D reconstruction, demonstrating the fastest on-the-fly training, superior representation quality, and real-time rendering capability. Project page: https://www.liuzhening.top/DASS
Figure 1: (Left) An overview of our three-stage pipeline, referred to as DASS. (Middle) Performance of 4D dynamic spatial reconstruction methods in terms of per-timestep training time and reconstruction quality (PSNR) on the N3DV dataset [26]. Online methods are represented by circles, while offline methods are indicated by triangles. The size of each marker is proportional to the rendering speed (FPS). Our method achieves the fastest online training speed, superior reconstruction quality, and real-time rendering capability. (Right) Visual comparisons of baseline methods, illustrating that our method preserves fine-grained details and recovers diverse dynamics.
@@ 1 Introduction
In recent years, Neural Radiance Field (NeRF) [32] has emerged as a promising approach for spatial representation from multi-view inputs, which optimizes neural networks to estimate color and density based on spatial positions and viewpoints. The extensions of NeRF to dynamic scene reconstruction have demonstrated remarkable effectiveness [26, 34, 36, 14, 16], yielding photo-realistic novel view synthesis. However, the efficiency of NeRF-based methods is severely hindered by their low rendering speed due to the dense queries of neural networks. To address this issue, 3D Gaussian Splatting (3DGS) [21] has been proposed as a solution to provide high-quality reconstruction and real-time rendering capabilities, leveraging its flexible point-based primitive design and tile-based differentiable rasterization. Subsequent efforts [19, 51, 27, 46, 3] have been dedicated to applying Gaussian Splatting for 4D dynamic reconstruction, with representative works integrating the time dimension into each Gaussian primitive [51] or learning spatio-temporal deformations [27, 46].
@@
Despite these advancements, most NeRF-based and 3DGS-based methods for 4D dynamic spatial reconstruction rely on full-length multi-view video inputs, which are non-causal in nature. This reliance limits their applicability in scenarios such as live streaming, where only per-timestep causal inputs are available, necessitating on-the-fly training. This problem can be formalized as the iterative reconstruction of 3D space at the current timestep, leveraging the reconstruction caches from the previous timestep and the multi-view inputs of the current timestep. The key challenges in this context are two-fold: (i) how to effectively model temporal variations in 3D space across consecutive timesteps and (ii) how to facilitate efficient optimization convergence from the previous timestep to the current one. Critical performance metrics for this problem include both the quality of novel view synthesis and the time efficiency.
@@
One intuitive solution is to optimize a new set of 3DGS primitives for each timestep independently. However, this strategy incurs substantial computational and storage costs, as it requires tuning and storing all 3DGS parameters for every timestep. A representative baseline, 3DGStream [37], addresses this issue by optimizing the transformation of Gaussian positions and rotation quaternions while adaptively densifying new Gaussians. Although this method achieves fast and high-quality reconstruction, it treats the entire scene uniformly and fails to distinguish between inherent dynamics and statics in the scene. In practice, dynamic and static components have different deformation characteristics. For instance, moving objects, such as humans or animals, display substantial dynamics, with the Gaussian properties like positions experiencing significant offsets. In contrast, static background and stationary objects show minimal movement, where Gaussians remain unchanged or undergo slight jitters. Consequently, uniformly modeling the transformation of all Gaussians serves as a sub-optimal strategy. Moreover, renewing the added Gaussian primitives for each timestep overlooks the temporal continuity inherent in the scene, leading to inefficiencies in both computation and representation.
@@
Based on these insights, we propose a dynamics-aware 3DGS streaming paradigm for on-the-fly 4D reconstruction, termed DASS, where the optimization of each timestep comprises three stages: inheritance, shift, and densification. Specifically, considering the temporal continuity, the newly densified Gaussians in the previous timestep are likely to persist in subsequent timesteps. Therefore, instead of renewing and optimizing added Gaussians for each timestep from scratch, we propose a selective inheritance mechanism to adaptively include a portion of the added Gaussians from the previous timestep using a learnable selection mask. Then, in the shift stage, we employ 2D dynamics-related priors, optical flow [48] and Gaussian segmentation [52] to calculate a per-Gaussian dynamics mask. Subsequently, we assign deformation layers to learn the offsets of dynamic and static Gaussians with different representation complexities, boosting a more dedicated optimization. In the densification stage, apart from the Gaussian offsets that present the deformations of existing objects, new Gaussian primitives are introduced to accommodate newly emerging objects. In this stage, both positional gradients and error maps from the shift stage serve as criteria for identifying regions that require densification. Iteratively, the inheritance stage in the subsequent timestep will process these added Gaussians, thereby mitigating errors in the shift stage and reducing the optimization burden in the densification stage. Our three-stage pipeline effectively captures dynamic spatial components and exploits the temporal correlation, achieving fast and high-fidelity on-the-fly 4D reconstruction. Our main contributions are summarized as follows:
@@
  • We propose a novel three-stage pipeline for 4D dynamic spatial reconstruction that supports on-the-fly training and per-timestep streaming. Our method builds on causal inputs and eliminates the need for full-length multi-view videos, thereby enhancing the practicability.
@@
  • Our approach seamlessly integrates three stages to optimize reconstruction quality. By selectively inheriting Gaussians from the preceding timestep, effectively distinguishing dynamic primitives to allocate optimization emphasis, and refining under-reconstructed areas using gradient information and optimization errors, our method ensures high-fidelity dynamic spatial reconstruction.
@@
  • Extensive experiments demonstrate the superiority of our method in multiple aspects, including the highest online training speed, superior reconstruction quality, and real-time rendering capability.
@@ 2.1 Neural Static Scene Representation
In recent years, reconstructing 3D representations from 2D plane visual inputs has experienced significant advancements, driven by the development of NeRF [32] and 3DGS [21]. NeRF-based methods represent spatial scenes by optimizing multi-layer perceptrons (MLPs) and generate novel views through volume rendering [18]. Subsequent research has enhanced both the training and rendering efficiency through grid-based designs [33, 49, 13, 14]. Nonetheless, NeRF-based approaches typically require dense ray tracing and struggle to fulfill high-speed rendering. Recently, 3DGS [21] has emerged to address these limitations by utilizing explicit unstructured scene representation while preserving point-based differentiable splatting rendering [57]. This approach achieves real-time rendering speed and photo-realistic quality. Based on these advancements, subsequent studies have further enhanced the representation efficiency [30, 7, 29, 10, 24, 55, 11, 53] and expanded applications in understanding and editing [52, 56, 6, 35, 43].
@@ 2.2 Neural Dynamic Scene Reconstruction
Extending static scene representation to dynamic FVV reconstruction remains a significant challenge, primarily due to the difficulties in modeling temporal correlations and variations in 3D space. To address this issue, several studies have extended NeRF to incorporate spatio-temporal structures [25, 5, 36, 14, 41], facilitating dynamic space reconstruction. Similarly, dynamic scene reconstruction methods based on Gaussian Splatting have been proposed, which can be categorized into deformation-based methods, 4D primitive-based methods, and iterative streaming methods. Deformation-based methods [46, 31, 50, 3] maintain 3D Gaussian representations and optimize neural deformable fields to capture temporal variations. 4D primitive-based methods [51, 27, 54] augment Gaussian primitives by integrating the temporal axis as an intrinsic property, thereby directly learning the spatio-temporal distributions. Note that both categories of methods rely on full-length multi-view videos for training, which limits their ability to support on-the-fly training and per-timestep streaming. In contrast, iterative streaming methods build upon previously converged representations and perform per-timestep optimization for each timestep’s multi-view inputs, thereby addressing the above limitations. However, this area remains largely underexplored. Our work aims to develop an iterative streaming pipeline that accelerates per-timestep optimization convergence and enhances the reconstruction quality.
The work most related to ours is 3DGStream [37]. It proposes a two-stage optimization process for each timestep, where the first stage optimizes a grid-based MLP [33] for Gaussian property transformation and the second stage wisely adds new Gaussians based on positional gradients. However, 3DGStream does not consider the intrinsic difference between dynamic and static features in the scene and discards the added primitives at each timestep, thereby limiting its ability to leverage temporal continuity across timesteps. Concurrently, other studies on iterative 4D reconstruction focus on different aspects, such as reducing streaming storage overhead on mobile devices [42] and exploring integration into communication systems [28, 38]. In contrast, our work focuses on achieving fast convergence in per-timestep optimization while maintaining high reconstruction quality, which is orthogonal to the contributions of aforementioned studies.
@@
Figure 2: Overview of our proposed DASS framework. The selective inheritance stage (Green) exploits the temporal continuity and adaptively preserves Gaussians from the previous timestep. The dynamics-aware shift stage (Blue) distinguishes the dynamic and static elements and optimizes the deformations. The error-guided densification stage (Yellow) detects and densifies the areas with weak reconstruction based on positional gradients and distortions. Variables highlighted in red represent learnable parameters in each stage, whose training is significantly lightweight compared to tuning all Gaussian parameters.
@@ 3.1 Overview
Our work addresses the problem of Gaussian Splatting streaming, which is formalized as an iterative 4D dynamic spatial reconstruction process, enabling on-the-fly training and per-timestep streaming. At each timestep $t$, the optimization begins with the representation from the previous timestep, denoted by $\bm{\mathcal{G}}_{t-1}$, and derives the current representation $\bm{\mathcal{G}}_{t}$, using multi-view inputs from $C$ viewpoints.
Our streaming pipeline, referred to as DASS, represents the 3D space at each timestep $t$ using two sets of Gaussian primitives: quantity-consistent base Gaussians and temporally quantity-varied densified Gaussians, i.e., $\bm{\mathcal{G}}_{t}:=\{\bm{\mathcal{G}}_{t}^{\textrm{base}},\bm{\mathcal{G}}_{t}^{\textrm{var}}\}$. The base Gaussians $\bm{\mathcal{G}}^{\textrm{base}}$, initialized from the zero timestep representation $\bm{\mathcal{G}}_{0}$, constitute the fundamental elements shared across all timesteps and maintain a fixed number of $N$ Gaussian primitives throughout the streaming process. In contrast, the number of Gaussians in $\bm{\mathcal{G}}^{\textrm{var}}$ varies across timesteps, capturing per-timestep unique emerging objects that are not represented in $\bm{\mathcal{G}}^{\textrm{base}}$.
@@
Fig. 2 depicts an overview of our framework, where the optimization for each timestep comprises three stages: inheritance, shift, and densification. The selective inheritance stage (Sec. 3.2) preserves temporal consistency by incorporating all base Gaussians $\bm{\mathcal{G}}_{t-1}^{\textrm{base}}$ from the previous timestep, while selectively inheriting the temporally-varied Gaussians $\bm{\mathcal{G}}_{t-1}^{\textrm{var}}$ using a per-Gaussian learnable mask. This mask effectively eliminates redundant Gaussians while retaining those essential for the current timestep. Then, the shift stage (Sec. 3.3) transforms the inherited Gaussians to align with the current timestep, capturing the movements of existing elements. In this stage, a Gaussian-level dynamics mask is estimated to classify the Gaussians into dynamic and static groups, which are then processed by two hash-encoding MLPs with different complexities. This shift stage efficiently manages the movements and rotations of Gaussians, accommodating changes from the previous timestep. In the densification stage (Sec. 3.4), emerging objects, such as coffee being poured from a cup or flames emanating from an oven, are identified by analyzing positional gradients and distortions resulting from the shift stage. This stage densifies a limited number of additional Gaussians to accurately represent these emerging objects. Our proposed design facilitates fast per-timestep convergence and achieves high-fidelity novel view synthesis. The details of each stage are elaborated in the following subsections.